# Task Aware Dreamer for Task Generalization in Reinforcement Learning

### v4

@@ Abstract
A long-standing goal of reinforcement learning is to acquire agents that can learn on various training tasks and generalize well on unseen tasks that may share a similar dynamic but with different reward functions. The ability to generalize across tasks is significant for real-world applications where the robot needs to adapt to varying reward mechanisms with the same embodiment. In this work, we first indicate that training general world models can utilize similar structures in these tasks and benefit training more generalizable agents. Extending world models into the task generalization setting, we introduce a novel method named Task Aware Dreamer (TAD), which integrates reward-informed features to identify consistent latent characteristics across tasks. Within TAD, we compute the variational lower bound of sample data log-likelihood, which introduces a new term designed to differentiate tasks using their states, as the optimization objective of our reward-informed world models. To demonstrate the advantages of the reward-informed policy utilized in TAD for handling the task distribution, we introduce a novel metric named Task Distribution Relevance (TDR) which quantitatively measures the relevance of different tasks. For tasks exhibiting a high TDR, i.e., these tasks differ significantly, we illustrate that Markovian policies struggle to distinguish them, thus it is necessary to utilize reward-informed policies in TAD. Extensive experiments in both image-based and state-based settings show that TAD can significantly improve the performance of handling different tasks meanwhile, especially for those with high TDR, and display a strong generalization ability to unseen tasks.
Figure 1: An overview. Given a task distribution, we train the agent in training tasks and hope it to zero-shot generalize to test tasks. For improving the generalization, we propose TAD, which utilizes $\Pi_{3}$ to encode all historical information for inferring the current task and novel reward-informed world models for capturing invariant latent features.
@@ 1 Introduction
Deep Reinforcement Learning (DRL) has demonstrated significant advancements in diverse fields (Mnih et al. 2016; Silver et al. 2016), and a key factor in these achievements is an agent’s proficiency in assimilating lessons from special training tasks. This tendency towards specialization harms the wider real-world application of DRL, where broad generalization across various tasks is necessary. One primary challenge in generalization is the correspondence in the underlying dynamics of different tasks, which exhibit distinct reward structures. For example, we may need to control the same robot to handle different tasks, which can be rewarded for speed or caution in a navigating robot. Developing generalizable agents that recognize and handle these subtle variations is still an area of keen interest and exploration in DRL.
For developing such generalizable agents, a promising pathway is to train general world models (Ha and Schmidhuber 2018; Hafner et al. 2019b, a) that help the agent understand the world and make decisions. In this work, we first provide insight that general world models benefit improving the sample efficiency in handling the task distribution, especially by utilizing similar dynamic structures of trajectories sampled from different tasks (Theorem 1). Consequently, we propose a novel framework named Task Aware Dreamer (TAD) to improve the generalization ability of agents via general world models. As current world models are primarily designed for the single-task setting (left of Fig. 2), we consider the corresponding probabilistic graphical model for the task-distribution setting (right of Fig. 2) and propose novel reward-informed world models that can capture invariant structures across tasks in the distribution. Then we compute the variational lower bound of the data log-likelihood as the primary training objective for optimizing the reward-informed world models, which incorporate a novel task context term that categorizes tasks based on their states. In practice, we implement this optimization objective via two alternative methods: cross-entropy (TAD-CE) and supervised-contrastive (TAD-SC).
Figure 2: Probabilistic graphical model designs for the single-task setting (left) and the task-distribution setting (right). The latter inspires the design of reward-informed world models. Solid and dashed lines represent the generative process and the inference model, respectively.
Moreover, we theoretically explain TAD’s components, like reward-informed policies and task optimization terms, are effective for handling task distribution. About the policy hypothesis, although several meta RL methods (Zintgraf et al. 2019; Rimon et al. 2024) have chosen the reward-informed policy hypothesis $\Pi_{3}$, the relationship between the task distribution and the expressiveness of the policy hypothesis set is still unclear. In other words, why the commonly adopted hypothesis set $\Pi_{1}$ of Markovian policies and the set $\Pi_{2}$ of policies encoding historical states and actions are not suitable for task distribution? To answer this question, we propose a novel metric of Task Distribution Relevance (TDR), encapsulating the relevance of different tasks within the distribution through their optimal Q functions. We then prove that both $\Pi_{1}$ and $\Pi_{2}$ are sub-optimal under the task-distribution setting. This sub-optimality is related to TDR (Theorem 3), i.e., for task distributions with high TDR, the performance of these two policy hypotheses might degenerate significantly, a phenomenon we also demonstrate in experiments (Sec. 5.2). This result explains why TAD and previous meta RL methods choose $\Pi_{3}$ for handling the task distribution. Besides the policy chosen, we also discuss our task optimization term in TAD and prove that it can effectively reduce the gap between the policy return and the optimal return (Theorem 4).
We evaluate the task generalization ability of TAD in extensive experiments, including DeepMind control suite (Tassa et al. 2018) and MuJoCo (Todorov, Erez, and Tassa 2012), which are image-based and state-based respectively. Agents are trained on various tasks and evaluated on unseen tasks. Results corroborate our analyses, indicating that $\Pi_{1}$ and $\Pi_{2}$ falter in managing task distributions characterized by a high TDR. Contrastingly, TAD excels in simultaneously managing varied tasks and outperforms all baselines, including SOTA model-based meta RL method MAMBA (Rimon et al. 2024). Additionally, our ablation studies highlight TAD’s versatility, showing its prowess in dynamic generalization and handling cross-embodiment tasks. Overall, our contributions include:
@@
  • We present theoretical insights that general world models can utilize similar structures across tasks and improve the sample efficiency for task generalization (Sec. 4.1).
  • Our TAD extends world models for the task distribution with the corresponding variational lower bound (Sec. 4.2).
@@
  • We theoretically analyze TAD’s components are effective for task generalization, with a novel metric TDR to quantify the distribution relevance (Sec. 4.3).
@@
  • Extensive experiments show that TAD can outperform various SOTA baselines and exhibit better generalization flexibility over image-based and state-based settings (Sec. 5).
@@ 2 Related Work
Current RL methods always struggle to generalize to new tasks (Song et al. 2019). Prior works have studied an array of training strategies, like loss regularization (Cobbe et al. 2019; Wang et al. 2020), successor representation (Touati and Ollivier 2021; Touati, Rapin, and Ollivier 2022), network architecture design (Lee et al. 2019; Raileanu and Fergus 2021), lifelong learning (Chen and Liu 2018; Mendez, van Seijen, and Eaton 2022), data augmentation (Raileanu et al. 2021; Hansen and Wang 2021), etc. Besides these, some works investigate the connection between policy generalization and the distribution of those environments. Ghosh et al. demonstrates that generalizing to unseen environments introduces partial observability, thereby rendering deterministic Markovian policies sub-optimal. Also, some studies (Lee et al. 2020; Ghosh et al. 2021) experimentally indicate that stochastic or non-Markovian policies can improve the generalization ability. However, the expressive abilities of differing hypothesis sets and their connection to the environment distribution, which are significant for developing more generalizable agents, remain understudied.
@@
These two topics are closely related to generalization in RL. Multi-task RL (Yang et al. 2020; Sodhani, Zhang, and Pineau 2021; Lee et al. 2022; Xu et al. 2022a) primarily aims to excel across all training tasks but is difficult to zero-shot generalize to unseen tasks. For boosting the generalization, Meta RL seeks to enable the trained agents to adapt to new tasks with few episodes, including gradient-based (Finn, Abbeel, and Levine 2017) and context-based methods (Duan et al. 2016; Rakelly et al. 2019). There are also some model-based methods (Nagabandi et al. 2018; Rimon et al. 2024) utilize learned models to boost the sample efficiency. Though some context-based methods like VariBAD (Zintgraf et al. 2019) show zero-shot generalization ability, it is still significant to directly analyze the generalization in RL and design corresponding algorithms.
@@
World models (Ha and Schmidhuber 2018) aims to better learn environmental representations, which has potential advantages for generalization as it can capture invariant features across tasks. Classical methods utilize the Recurrent State Space Model (RSSM) (Hafner et al. 2019b) for planning (Hafner et al. 2019b) and policy learning (Hafner et al. 2019a, 2020, 2023). Subsequent research explored reconstruction-free world models (Deng, Jang, and Ahn 2022), temporal predictive coding (Nguyen et al. 2021), cooperative reconstruction (Fu et al. 2021), and Denoised MDP (Wang et al. 2022) for more effective task-relevant information encoding. World models are also utilized to extract environmental invariant features by learning from videos (Seo et al. 2022) or exploration (Sekar et al. 2020; Xu et al. 2022b), and then fine-tuning to new tasks. However, most world models are designed for the single-task setting and struggle to manage multiple tasks without fine-tuning, limiting their effectiveness for zero-shot generalization to unseen tasks.
@@ 3 Preliminary
We consider the setting with a task distribution $\mathcal{T}$ of Partially observable Markov decision processes (POMDPs), where different tasks own the same dynamic and different rewards. Formally, each POMDP $\mathcal{M}\sim\mathcal{T}$ can be represented as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\Omega,\mathcal{O})$. Here $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively. For $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, $\mathcal{P}(\cdot|s,a)$ is the state transition probability that is Markovian, $\mathcal{R}_{\mathcal{M}}(s,a)$ is its reward function. The state is unobserved for the agent, which can only access the observation from the observation space $\Omega$ calculated by the observation function $\mathcal{O}(\cdot|s)$.
Following previous meta RL and generalization methods, we consider policies that encoder all historical information (we prove its necessity in Sec. 4.3). Formally, at each timestep $t$, the agent with the policy $\pi$ will use the whole history trajectory $(o_{0},a_{0},r_{0},o_{1},...,o_{t})$ to sample action $a_{t}$, arrive at the next state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$, and get the current reward $r_{t}=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$. The performance of policy $\pi$ in $\mathcal{M}$ is defined as the expected discounted return: $J_{\mathcal{M}}(\pi)=\mathbb{E}_{\tau\sim\pi}\left[R(\tau)\triangleq\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$. Our objective is to maximize the return over $\mathcal{T}$, i.e., $\max_{\pi}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]$.
In practice, given the task distribution $\mathcal{T}$, we sample $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ for optimizing the agent, i.e., maximizing $\frac{1}{M}\sum_{m=1}^{M}J_{\mathcal{M}_{m}}(\pi)$. In the testing stage, we will sample $N$ unseen test tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$ to evaluate its generalization ability, i.e., evaluating $\frac{1}{N}\sum_{n=1}^{N}J_{\mathcal{M}_{M+n}}(\pi)$.
@@ 4 Task Aware Dreamer
In this section, we first demonstrate that general world models benefit task generalization. Then we introduce our reward-informed world models and propose a novel framework of Task Aware Dreamer (TAD) for handling task generalization. Finally, we provide theoretical analyses of designs in TAD.
@@ 4.1 Reward-Informed World Models
Our first observation is that general world models are effective in narrowing down the hypothesis space of the optimal Q function when handling task generalization:
Theorem 1 (Proof in Appendix A.1).
@@
Set $\mathcal{Q}$ as the space of observation-action Q functions. Given $M$ tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ and corresponding dataset $\mathcal{D}_{m}=\{(o_{t}^{m},a_{t}^{m},r_{t}^{m},o_{t+1}^{m})\}$, we set the product space $\mathcal{H}=\mathcal{Q}^{M}$ composed of M spaces, i.e., $\forall\{q_{m}\}_{m=1}^{M}\in\mathcal{H}$, $q_{m}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ belongs to $\mathcal{Q}$. Considering the following three hypothesis classes $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3}\subseteq\mathcal{H}$:
$\begin{split}\mathcal{H}_{1}=&\{(q_{m})_{m=1}^{M}|q_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m}+\gamma\max_{a^{\prime}}q_{m}(o_{t+1}^{m},a^{\prime})\}\\ \mathcal{H}_{2}=&\{(q_{m})_{m=1}^{M}|\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall o,a\}\\ \mathcal{H}_{3}=&\{(q_{m})_{m=1}^{M}|\exists p:p(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},\\ &q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p(o,a),a^{\prime}),\forall o,a\}.\end{split}$
Then we have $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$.
@@
Here $\mathcal{H}_{1}$ own Q functions satisfying optimal Bellman equation with data from $\mathcal{D}_{m}$, while $\mathcal{H}_{2}$ and $\mathcal{H}_{3}$ own Q functions satisfying optimal Bellman function with data from world models, which are trained with the single dataset and all datasets, respectively. Consequently, $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$ utilizes the generalization ability of world models, which extends the results in the fixed-task setting (Young et al. 2023). Moreover, $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$ illustrate similar dynamic structures across different tasks benefit learning general world models. Consequently, training general world models can narrow down the hypothesis of possible Q function and benefit agent training.
@@
As existing world models (Hafner et al. 2019b, a) are mainly designed for the single-task setting, we propose reward-informed world models for the task-distribution setting. We first analyze the probabilistic graphical model of the task-distribution setting shown in Fig. 2, where rewards do not only rely on the previous states and actions but also the current task. Thus we can calculate the joint distribution as
$p(s_{1:T},o_{1:T},r_{1:T},a_{1:T-1},\mathcal{M})$ (1)
$=$ $p(\mathcal{M})\prod_{t=1}p(s_{t+1}|s_{t},a_{t})p(o_{t}|s_{t})p(r_{t}|s_{t},\mathcal{M}),$
@@
and we choose the reward-informed inference model (the necessity of reward-informed policy in task generalization is discussed in Sec. 4.3) to approximate state posteriors as
@@
Based on this inference model, we can construct the variational lower bound of the log-likelihood on the data as
$\begin{split}&\ln p(o_{1:T},r_{1:T},\mathcal{M}|a_{1:T})\\ \geq&\sum_{t=1}^{T}\mathbb{E}_{q(s_{t}|o_{\leq t},a_{<t},r_{<t})}\left[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})\right.\\ -&\left.\text{KL}\left(q(s_{t}|o_{\leq t},r_{<t},a_{<t})\|p(s_{t}|s_{t-1},a_{t-1})\right)\right]\\ +&\mathbb{E}_{q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})}\left[\ln p(\mathcal{M}|s_{1:T})\right].\end{split}$ (3)
This result is a general form of the single-task setting (Hafner et al. 2019b) and its derivation is in Appendix A.2. The first two terms in Eq.(LABEL:eq_elbo) are for reconstructing observations, predicting rewards, and inferring states, which are similar to the single-task setting. The last novel term is dedicated to predicting the current task from historical information, which is beneficial for improving the generalization as it encourages inferring the current task context. Extending RSSM via Eq.(LABEL:eq_elbo), our reward-informed world models consist of:
Deterministic state model: $h_{t}=f(h_{t-1},s_{t-1},a_{t-1},r_{t-1}),$
@@
Here hidden state $h_{t}$ encodes historical states, actions, and rewards, by using gated recurrent unit (GRU) (Chung et al. 2014) as $f(\cdot)=\text{GRU}(\cdot)$. Then the transition model, observation model, reward model, and task model further predict state, observation, reward, and task context respectively.

### v5

@@ Abstract
A long-standing goal of reinforcement learning is to achieve agents that can learn on various training tasks and generalize well on unseen tasks with similar transition dynamics but different reward functions. For instance, a household robot needs to use the same embodiment to complete various tasks characterized by distinct reward functions. To develop agents capable of handling this challenge, we construct Reward-Informed World Models (RIWM) that learns task-specific features while utilizing similar transition dynamics of trajectories collected from different tasks. The world model is optimized with a novel algorithm named Task Aware Dreamer (TAD), which employs a task context term to guide the world model to differentiate tasks and thus enhance task generalization. We further provide theoretical analyses to show that TAD’s key components are essential for handling task generalization. To establish the necessity of the policy hypothesis set utilized in TAD, we introduce a novel metric named Task Distribution Relevance (TDR), which quantitatively measures the relevance of different tasks. Tasks with high TAD often have vastly different optimal polices and thus are impossible to solve with traditional Markovian policy hypothesis sets, thereby necessitating the reward-informed policy used in TAD. Extensive experiments in both image-based and state-based benchmarks show that TAD can significantly improve the performance of handling various training tasks, especially for those with high TDR, and exhibits strong zero-shot generalization ability to unseen tasks. We also conduct ablation studies that demonstrate TAD’s potential to handle transition variations in cross-embodiment scenarios.
keywords:
reinforcement learning, task generalization, world models, model-based reinforcement learning, deep learning
@@ 1 Introduction
Although Deep Reinforcement Learning (DRL) agents have demonstrated significant advancements in various fields [1, 2], their achievements are usually limited to specific training tasks. This tendency toward specialization harms the real-world applicability of DRL agents, where robust generalization across diverse tasks is necessary. For example, a household robot needs to handle different tasks that share the same underlying transition dynamics (i.e., the same embodiment and physical environment) but are characterized by distinct reward functions. Developing generalizable agents that can recognize and handle these subtle variations remains an active and important area of research in DRL.
Compared to the traditional single-task setting, we formalize the problem as the task-distribution setting (Sec. 3), where each training/testing task is sampled from a task distribution $\mathcal{T}$. To develop agents capable of handling this situation, we show that the world model [3, 4, 5] is a promising pathway to address the challenges posed by the variability of tasks. In this work, we first demonstrate that world models can enhance the sample efficiency when handling the task-distribution setting, particularly by exploiting shared dynamic structures of trajectories sampled from different tasks (Theorem 4.1). While current world models are primarily tailored for the single-task setting (left of Fig. 1b), we extend the probabilistic graphical model to handle the task-distribution setting (right of Fig. 1b) by introducing the task random variable $\mathcal{M}$. Building upon this framework, we extend the traditional world [3, 4, 5] models to our Reward-Informed World Models (RIWM, Eq. (LABEL:eq_world_model)). RIWM incorporates historical reward signals to optimize latent states and integrates a task model to infer the task context of the trajectory. We further propose a novel algorithm named Task Aware Dreamer (TAD) to optimize the aforementioned RIWM by maximizing the variational evidence lower bound. The derived task context term encourages the world model to learn compact representations that effectively distinguish among tasks, thereby enhancing the agents’ ability to generalize across tasks. In practice, we implement this objective using two alternative approaches: a cross-entropy–based method (TAD-CE) and a supervised contrastive method (TAD-SC).
(a) Overview of this work.
(b) Probabilistic graphical model designs.
Figure 1: (a) An overview. Given a task distribution, the agent is trained on the training tasks and is expected to generalize to the test tasks in a zero-shot manner. Here $\Pi_{1}$ represents the set of Markovian policies, while $\Pi_{2}$ denotes all policies that include historical state and action information. To improve generalization ability, we propose TAD, which utilizes $\Pi_{3}$ to encode all historical information for inferring the current task and introduces novel Reward-Informed World Models to capture task-specific latent features. (b) Probabilistic graphical model designed for the single-task setting (left) and the task-distribution setting (right). The latter illustrates the design of Reward-Informed World Models, where solid and dashed lines denote the generative process and the inference model, respectively.
To validate the effectiveness of the proposed method, we theoretically demonstrate that the key innovations of TAD are essential for addressing task generalization. Specifically, we first establish the necessity of adopting the reward-informed policy hypothesis $\Pi_{3}$ in the task-distribution setting. Although several meta RL methods [6, 7] practically choose $\Pi_{3}$, the relationship between the task distribution and the expressiveness of the policy hypothesis set remains insufficiently understood. In particular, we address the following question: Why are the commonly adopted hypothesis set $\Pi_{1}$ of Markovian policies and set $\Pi_{2}$ of policies that encode historical states and actions inadequate for the task-distribution setting?
To answer this question, we propose a novel metric named Task Distribution Relevance (TDR), which encapsulates the relevance among tasks within the distribution through their optimal Q-functions. We then prove that both $\Pi_{1}$ and $\Pi_{2}$ are sub-optimal in the task-distribution setting. This sub-optimality is characterized by TDR (Theorem 4.4). Specifically, for task distributions with high TDR, the performance of these two policy hypothesis sets can degrade significantly, a phenomenon we also empirically demonstrate in experiments (Sec . 5.2). This result provides a theoretical justification for the adoption of $\Pi_{3}$ in TAD as well as in prior meta-RL methods when addressing task-distribution settings. Beyond the choice of policy hypothesis set, we further analyze the task context term in TAD and prove that it effectively reduces the gap between the policy return and the optimal return (Theorem 4.5).
We conduct extensive experiments to evaluate TAD’s ability to enhance task generalization, using commonly adopted benchmarks from the DeepMind Control Suite [8] and MuJoCo [9]. Both training and testing tasks are drawn from a task distribution $\mathcal{T}$, which requires the agents to simultaneously learn from multiple training tasks and generalize to potentially unseen test tasks. The experimental results corroborate our theoretical analysis, showing that policies from $\Pi_{1}$ and $\Pi_{2}$ struggle to handle tasks from distributions with a high TDR. Conversely, TAD not only effectively handles diverse training tasks but also exhibits better zero-shot generalization performance than the few-shot adaptation performance of the SOTA meta-RL method MAMBA [7]. Moreover, our ablation studies demonstrate the versatility of TAD, highlighting its potential for dynamics generalization and handling cross-embodiment task settings.
Overall, our contributions are summarized as follows:
@@
  • We provide theoretical insights showing that world models can exploit shared transition dynamics across tasks to improve sample efficiency for task generalization (Sec. 4.1).
  • We extend world models to the task-distribution setting by proposing TAD, which introduces a task context term that encourages representations to capture task-relevant information (Sec. 4.2).
@@
  • We theoretically demonstrate the effectiveness of TAD’s components for task generalization and introduce a novel metric named TDR to quantify task distribution relevance (Sec. 4.3).
@@
  • Extensive experiments show that TAD outperforms a range of SOTA baselines and exhibits strong generalization flexibility across both image-based and state-based benchmarks (Sec. 5).
@@ 2 Related work
Current RL agents often struggle to generalize to new tasks [10]. Prior work has explored a wide range of training strategies to address this challenge, including loss regularization [11, 12], successor representations [13, 14], network architecture design [15, 16], data augmentation [17, 18], leveraging trajectory similarities [19], and contrastive objectives [20]. Beyond these approaches, several studies have investigated the relationship between policy generalization and the distribution over environments. In particular, LEEP [21] demonstrates that generalizing to unseen environments introduces partial observability, which makes deterministic Markovian policies suboptimal. Moreover, empirical evidence from prior works [22, 21] suggests that stochastic or non-Markovian policies can enhance generalization ability. Despite these insights, a theoretical gap remains regarding the expressiveness of different hypothesis sets and their relationship to the environment distribution, both of which are crucial for developing more generalizable agents.
@@
These two topics are closely related to generalization in RL. Multi-task RL [23, 24, 25, 26] primarily focuses on excelling across all training tasks, but often struggles to generalize to previously unseen tasks. In contrast, Meta RL aims to equip agents with the ability to rapidly adapt to new tasks with a few episodes, through approaches such as gradient-based [27] and context-based methods [28, 29]. In addition, several model-based approaches [30, 7] leverage learned environment models to improve sample efficiency. Although certain context-based methods, such as VariBAD [31], demonstrate zero-shot generalization capabilities, it remains important to directly study the mechanisms of generalization in RL and to develop algorithms specifically tailored to this objective.
@@
World models [3] aim to learn rich representations of the environment, offering potential advantages for generalization by capturing task-invariant features. Classical approaches typically adopt the Recurrent State Space Model (RSSM) [4] for planning [4] and policy learning [5, 32, 33]. Subsequent work has proposed reconstruction-free world models [34], temporal predictive coding [35], cooperative reconstruction [36], and Denoised MDPs [37] to more effectively encode task-relevant information. World models have also been leveraged to extract environment-invariant features through videos [38] or unsupervised exploration [39, 40], followed by fine-tuning on downstream tasks. Despite these advances, most existing world models are designed for single-task settings and require task-specific fine-tuning to handle multiple tasks, which limits their ability to achieve zero-shot generalization to unseen tasks.
@@ 3 Preliminary
We consider a setting with a task distribution $\mathcal{T}$ over Partially Observable Markov Decision Processes (POMDPs), where different tasks share identical dynamics but differ in their reward functions. Formally, each POMDP $\mathcal{M}\sim\mathcal{T}$ is defined as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\Omega,\mathcal{O})$. Here, $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, respectively. For $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, $\mathcal{P}(\cdot|s,a)$ specifies the Markovian state transition dynamics, while $\mathcal{R}_{\mathcal{M}}(s,a)$ denotes the task-specific reward function. The underlying state $s$ is not directly observable by the agent. Instead, the agent receives observations from the observation space $\Omega$ according to the observation function $\mathcal{O}(\cdot\mid s)$.
Following prior work on meta-RL and generalization, we consider policies that encode all historical information (we prove the necessity of this assumption in Sec. 4.3). Formally, at each time step $t$, the agent with the policy $\pi$ will utilize the whole historical trajectory $(o_{0},a_{0},r_{0},o_{1},...,o_{t})$ to sample action $a_{t}$, arrive at the next state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$, and receive current reward $r_{t}=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$. The performance of policy $\pi$ on task $\mathcal{M}$ is measured by the expected discounted return $J_{\mathcal{M}}(\pi)=\mathbb{E}_{\tau\sim\pi}\left[R(\tau)\triangleq\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$. Our objective is to learn a policy that maximizes the expected return over the task distribution $\mathcal{T}$: $\max_{\pi}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]$.
In practice, given the task distribution $\mathcal{T}$, we sample $M$ training tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ to optimize the agent by maximizing the empirical objective $\frac{1}{M}\sum_{m=1}^{M}J_{\mathcal{M}_{m}}(\pi)$. At test time, we sample $N$ previously unseen tasks $\{\mathcal{M}_{M+n}\}_{n=1}^{N}$ to evaluate the agent’s generalization performance, measured by $\frac{1}{N}\sum_{n=1}^{N}J_{\mathcal{M}_{M+n}}(\pi)$.
@@ 4 Task aware dreamer
In this section, we begin by establishing that world models are beneficial for task generalization. We then introduce Reward-Informed World Models (RIWM) and propose Task-Aware Dreamer (TAD) as a novel approach to promote generalization. Finally, we provide theoretical analyses demonstrating that TAD can effectively enhance performance across various tasks.
@@ 4.1 Reward-informed world models
Our first observation is that world models are effective in narrowing the hypothesis space of the optimal Q-function when handling task generalization:
Theorem 4.1 (Proof in Appendix A.1).
@@
Set $\mathcal{Q}$ denotes the space of observation-action Q-functions. Given $M$ tasks $\{\mathcal{M}_{m}\}_{m=1}^{M}$ and corresponding dataset $\mathcal{D}_{m}=\{(o_{t}^{m},a_{t}^{m},r_{t}^{m},o_{t+1}^{m})\}$, we use $\mathcal{H}=\mathcal{Q}^{M}$ to represent the product space of $M$ Q-function spaces, i.e., $\forall\{q_{m}\}_{m=1}^{M}\in\mathcal{H}$, $q_{m}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ belongs to $\mathcal{Q}$. Considering the following three hypothesis sets $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3}\subseteq\mathcal{H}$:
$\begin{split}\mathcal{H}_{1}=&\{(q_{m})_{m=1}^{M}|q_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m}+\gamma\max_{a^{\prime}}q_{m}(o_{t+1}^{m},a^{\prime})\}\\ \mathcal{H}_{2}=&\{(q_{m})_{m=1}^{M}|\exists(p_{m})_{m=1}^{M}:p_{m}(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p_{m}(o,a),a^{\prime}),\forall o,a\}\\ \mathcal{H}_{3}=&\{(q_{m})_{m=1}^{M}|\exists p:p(o_{t}^{m},a_{t}^{m})=o_{t+1}^{m},\\ &\exists(r_{m})_{m=1}^{M}:r_{m}(o_{t}^{m},a_{t}^{m})=r_{t}^{m},q_{m}(o,a)=r_{m}(o,a)+\gamma\max_{a^{\prime}}q_{m}(p(o,a),a^{\prime}),\forall o,a\}.\end{split}$
The hypothesis sets satisfy $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$.
@@
Here, $\mathcal{H}_{1}$ contains Q-functions that satisfy the optimal Bellman equation with data from $\mathcal{D}_{m}$. In contrast, $\mathcal{H}_{2}$ and $\mathcal{H}_{3}$ contain Q-functions that satisfy the optimal Bellman equation using data generated by world models trained on a single dataset and on all datasets, respectively. Consequently, $\mathcal{H}_{2}\subseteq\mathcal{H}_{1}$ utilizes the generalization ability of world models, extending prior results in the fixed-task setting [41]. Moreover, $\mathcal{H}_{3}\subseteq\mathcal{H}_{2}$ suggests that shared dynamic structures across tasks benefit learning world models. Overall, Theorem 4.1 indicates that training world models narrows the hypothesis space of plausible Q-functions, thereby facilitating agent learning and improving generalization.
@@
As existing world models [4, 5] are primarily designed for the single-task setting, we propose Reward-Informed World Models (RIWM) for the task-distribution setting. We begin by analyzing the probabilistic graphical model of this setting, illustrated in Fig. 1b, where the rewards depend not only on previous states and actions but also on the current task. Accordingly, the joint distribution can be expressed as:
$p(s_{1:T},o_{1:T},r_{1:T},a_{1:T-1},\mathcal{M})=p(\mathcal{M})\prod_{t=1}p(s_{t+1}|s_{t},a_{t})p(o_{t}|s_{t})p(r_{t}|s_{t},\mathcal{M}),$ (1)
@@
and we choose to adopt the reward-informed inference model (the necessity of reward-informed policy in task generalization is discussed in Sec. 4.3) to approximate state posteriors as:
@@
Based on this inference model, we can construct the variational lower bound of the action-conditioned probability as:
$\begin{split}\ln p(o_{1:T},r_{1:T},\mathcal{M}|a_{1:T})\geq&\sum_{t=1}^{T}\mathbb{E}_{q(s_{t}|o_{\leq t},a_{<t},r_{<t})}\left[\ln p(o_{t},r_{t}|\mathcal{M},s_{t})-\mathrm{KL}\left(q(s_{t}|o_{\leq t},r_{<t},a_{<t})\|p(s_{t}|s_{t-1},a_{t-1})\right)\right]\\ &+\mathbb{E}_{q(s_{1:T}|o_{1:T},a_{1:T},r_{1:T})}\left[\ln p(\mathcal{M}|s_{1:T})\right].\end{split}$ (3)
This result is a general form of the single-task setting [4], with the full derivation in Appendix A.2. The first two terms in Eq. (3) correspond to reconstructing observations and predicting rewards, which are similar to the single-task setting. The final novel term is dedicated to predicting the current task from historical information, which enhances generalization by encouraging the model to infer the current task context. Building on RSSM via Eq. (3), our RIWM consists of:
Deterministic state model: $h_{t}=f(h_{t-1},s_{t-1},a_{t-1},r_{t-1}),$ (4)
@@
Here, the hidden state $h_{t}$ encodes historical states, actions, and rewards using a gated recurrent unit (GRU) [42], i.e., $f(\cdot)=\text{GRU}(\cdot)$. The transition, observation, reward, and task models then predict the next state, observation, reward, and task context, respectively.